Three times the complement of angle A is 40 degrees more than the supplement of angle A. Find the measure of angle A.
Answer by mananth(12270)

Three times the complement of angle A is 40 degrees more than the supplement of angle A. Find the measure of angle A.
..
let the angle be x
complement = 90-x
supplement = 180-x
....
180-x = 3(90-x) +40
180-x = 270-3x+40
3x-x=270-180+40
2x=130
x= 130/2
x= 65 deg.